# M3: A Modular M orld Model over Streams of Tokens

### v2

@@ 3 Experiments
To evaluate sample efficiency, we used benchmarks that measure performance within a fixed, limited environment interaction budget. The selected benchmarks also addressed two key research questions: (1) whether $\text{M}^{\text{3}}$ performs effectively in continuous environments and (2) whether it handles multi-modal observations successfully.
@@ 3.1 Experimental Setup
We evaluate $\text{M}^{\text{3}}$ on three sample-efficiency benchmarks of different observation and action modalities: Atari 100K (Kaiser et al., 2020), DeepMind Control Suite (DMC) Proprioception 500K (Tunyasuvunakool et al., 2020), and Craftax-1M (Matthews et al., 2024).
@@
The DeepMind Control Suite (DMC) is a set of continuous control tasks involving multiple agent embodiments ranging from simple single-joint models to complex humanoids. Here, we follow the subset of proprioception tasks used for the evaluation of DreamerV3 (Hafner et al., 2023), where observations and actions are continuous vectors. At each task, the agent’s interaction budget is limited to 500K steps.
@@
On Atari-100K, we compare $\text{M}^{\text{3}}$ against DreamerV3 (Hafner et al., 2023) and several methods restricted to image observations: TWM (Robine et al., 2023), STORM (Zhang et al., 2024), DIAMOND (Alonso et al., 2024), and REM (Cohen et al., 2024). On DMC, we compare exclusively with DreamerV3, currently the only planning-free world model method with published results on the 500K proprioception benchmark. On Craftax-1M, we compare against TWM (Dedieu et al., 2025), a concurrent work that proposes a Transformer based world model with a focus on the Craftax benchmark, and the baselines reported in the Craftax paper: Random Network Distillation (RND) (Burda et al., 2019), PPO (Schulman et al., 2017) with a recurrent neural network (PPO-RNN), and Exploration via Elliptical Episodic Bonuses (E3B) (Henaff et al., 2022). As Craftax is a recent benchmark, there are no other published results in existing world models literature. Following the standard practice in the literature, we exclude planning-based methods (Hansen et al., 2024; Wang et al., 2024), as planning is an orthogonal component that operates on any given model, typically incurring significant computational overhead. Here, our focus is on the world model component.
@@
For Atari, we report human-normalized scores (HNS) $\frac{\text{agent\_score}-\text{random\_score}}{\text{human\_score}-\text{random\_score}}$ (Mnih et al., 2015). Following the protocol of (Agarwal et al., 2021) and using their toolkit, we report the mean, median, interquantile mean (IQM), and optimality gap metrics with 95% stratified bootstrap confidence intervals. For DMC and Craftax, we report the raw agent returns. We use 5 random seeds per environment. In each experiment, final performance is evaluated using 100 test episodes at the end of training and the mean score is reported.
@@ 3.2 Results
$\text{M}^{\text{3}}$ achieves state-of-the-art performance across all three benchmarks (Figure 1). On Atari 100k, $\text{M}^{\text{3}}$ outperforms all baselines across all key metrics (Figure 4). Notably, $\text{M}^{\text{3}}$ is the first planning-free world model to reach a humen-level IQM and median scores. In addition, it achieves superhuman performance on 13 out of 26 games, surpassing all baselines. These results highlight $\text{M}^{\text{3}}$’s effectiveness in sample-efficient learning and its robustness across diverse tasks in the Atari 100k benchmark.
Is $\text{M}^{\text{3}}$ effective in continuous environments?
Figure 4 provides compelling evidence that token-based architectures can indeed excel in continuous domains: $\text{M}^{\text{3}}$ achieves performance comparable to DreamerV3 across most tasks, and notably outperforms it on Cartpole Swingup Sparse and both Hopper tasks.
Is $\text{M}^{\text{3}}$ effective in environments with multi-modal observations?
Since modularity is at the core of $\text{M}^{\text{3}}$, we investigate its performance in Craftax, as it combines an image-like 2D grid map with a vector of features, involving multiple tokenizers ($\mathcal{V}$). Figure 5 shows that $\text{M}^{\text{3}}$ maintains sample-efficiency in this multi-modal environment as well, as it significantly outperforms all model-free baselines, including exploration-oriented ones. $\text{M}^{\text{3}}$ also attains state-of-the-art performance compared to concurrent world model methods (Figure 1). With 444 tokens per observation arranged into 147 sequences, even short trajectories in Craftax contain thousands of tokens, demonstrating $\text{M}^{\text{3}}$’s efficient handling of long sequences.
@@ 3.3 Ablation Studies
We ablate the intrinsic rewards, prioritized replay, and classification-based predictions to demonstrate their individual contributions to $\text{M}^{\text{3}}$’s performance. In each ablation, $\text{M}^{\text{3}}$ is modified so that only the component of interest is disabled. Due to limited computational resources, we consider a subset of 8 tasks for each of the Atari 100K and DMC benchmarks. Concretely, for Atari 100K, we used ”Assault”, ”Breakout”, ”Chopper Command”, ”Crazy Climber”, ”James bond”, ”Kangaroo”, ”Seaquest”, and ”Up’n Down”, in which significant improvements were observed. For DMC, we chose a subset that includes different embodiments: ”acrobot swingup”, ”cartpole swingup sparse”, ”cheetah run”, ”finger turn hard”, ”hopper stand”, ”pendulum swingup”, ”reacher hard”, and ”walker run”.
The results are presented in Figure 6. Although all components contributed to the final performance of $\text{M}^{\text{3}}$, the intrinsic rewards proved crucial for achieving competitive performance, particularly on DMC.
More broadly, the results in Figure 6 demonstrate that encouraging the controller to explore regions of high epistemic uncertainty through intrinsic rewards significantly improves its performance in world model agents, even in reward-rich environments. This observation is non-trivial in a sample-efficiency setting, as model-driven exploration costs the agent some amount of interaction budget, which could be otherwise used for controller-driven exploration during data collection. The latter type of exploration aims to collect new information about the true reward signal, which defines the task and its success metric. On the other hand, model-driven exploration may guide the controller towards environment regions that are irrelevant to the task at hand.
@@ 4 Related Work
Large-scale token-based sequence models for modeling and generating multi-modal trajectories of agent experience were proposed in (Lu et al., 2023, 2024; Reed et al., 2022; Schubert et al., 2023). In Gato (Reed et al., 2022) and TDM (Schubert et al., 2023), multi-modal inputs are first tokenized through predefined transformations, while in Unified IO (Lu et al., 2023, 2024) pretrained models are also used. The produced tokens are then embedded and processed by the sequence model, similarly to $\text{M}^{\text{3}}$. Importantly, these methods do not learn control through RL. Instead, they learn from expert data. In addition, these methods use large models with billions of parameters, large vocabulary sizes, and orders-of-magnitude more data and compute compared to sample efficient world models. Hence, it is unclear whether the design choices of these methods would be effective in an online sample-efficiency setting, where the data is non-stationary and strictly limited.
@@
Model-based RL agents where the policy is learned exclusively from simulated data generated by the learned world model were originally proposed by (Ha & Schmidhuber, 2018). Later, a line of works proposed the popular Dreamer algorithms (Hafner et al., 2020, 2021, 2023). To learn dynamics, Dreamer’s objective involves a Kullback–Leibler (KL) divergence term between its learned prior and posterior estimators, effectively coupling the optimization of the representation model and the recurrent neural network world model. This design leads to a complex monolithic system that poses challenges for development and scaling.
Following the success of the Transformer architecture (Vaswani et al., 2017) in language modeling (Brown et al., 2020), Transformer-based variants of Dreamer (Zhang et al., 2024; Robine et al., 2023) were proposed. In addition, token-based world models (TBWMs), which represent trajectories as language-like token sequences, were proposed (Micheli et al., 2023; Cohen et al., 2024). Notably, the evaluation of these methods is limited to the Atari 100K benchmark.
Recently, motivated by the success of diffusion generative models (Rombach et al., 2022), DIAMOND (Alonso et al., 2024), a diffusion-based world model agent was proposed. Although it produces visually compelling outputs, it is currently limited to visual environments.
(a) Atari 100K
(b) DeepMind Control Suite Proprioception 500K
Figure 6: Ablations results on the Atari-100K (top) and DeepMind Control Proprioception 500K (bottom) benchmarks. A subset of 8 games was used for each ablation.
Large Video World Models
Following recent advances in video generative modeling (Ho et al., 2022; Blattmann et al., 2023; Brooks et al., 2024), a recent body of work proposed large video world models (DeepMind, 2024; Agarwal et al., 2025; Decart et al., 2024; Valevski et al., 2024). These methods are trained offline on massive pre-collected data to predict future video frames. However, these methods do not involve control learning, and RL in particular.
@@ 5 Limitations and Future Work
Here, we briefly highlight several limitations of this work. First, although the feature quantization approach for tokenizing continuous vectors showed promise, it leads to excessive sequence lengths. We believe that more efficient solutions can be found for dealing with continuous inputs. Second, since rich multi-modal RL benchmarks are scarce, future work could wrap multiple single-modality environments into a unified multi-modal environment, which can be used to evaluate agents under different modality combinations as well as the agent’s ability to perform multiple independent tasks concurrently.
@@ 6 Conclusions
In this paper, we proposed a modular world model, $\text{M}^{\text{3}}$, for sample-efficient RL. $\text{M}^{\text{3}}$ extends token-based world model agents via a modular framework to support a wide range of environments and modalities. In addition, it combines multiple advances from existing literature to enhance performance. Through extensive empirical evaluation, $\text{M}^{\text{3}}$ exhibited state-of-the-art performance for planning-free world models on a diverse set of benchmarks, involving visual, continuous, and structured symbolic modalities. Notably, on the well-established Atari-100K benchmark, $\text{M}^{\text{3}}$ outperformed all baselines across all metrics.
7 Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ A.1 Hyperparameters
Due to limited computational resources, $\text{M}^{\text{3}}$ was not properly tuned. We believe that proper tuning could further improve $\text{M}^{\text{3}}$’s performance.
@@ A.2.1 Image Observations
Image observations are tokenized using a vector-quantized variational auto-encoder (VQ-VAE) (van den Oord et al., 2017; Esser et al., 2021). A VQ-VAE comprises a convolutional neural network (CNN) encoder, an embedding table $\mathbf{E}\in\mathbb{R}^{n\times d}$, and a CNN decoder. Here, the size of the embedding table $n$ determines the vocabulary size.
@@
$z=\arg\min_{i}\|\mathbf{h}-\mathbf{E}(i)\|$
where $\mathbf{E}(i)$ is the $i$-th row of $\mathbf{E}$. To reconstruct the original image, the decoder first maps $\mathbf{z}$ to their embeddings using $\mathbf{E}$. During training, the straight-through estimator (Bengio et al., 2013) is used for backpropagating the learning signal from the decoder to the encoder: $\hat{\mathbf{h}}=\mathbf{h}+\operatorname{sg}(\mathbf{E}_{z}-\mathbf{h})$. The architecture of the encoder and decoder models is presented in Table 8.
@@
$\mathcal{L}(\operatorname{enc},\operatorname{dec},\mathbf{E})=\|\mathbf{o}-\operatorname{dec}(z)\|_{2}^{2}+\|\operatorname{sg}(\operatorname{enc}(\mathbf{o}))-\mathbf{E}(z)\|_{2}^{2}+\|\operatorname{sg}(\mathbf{E}(z)-\operatorname{enc}(\mathbf{o})\|_{2}^{2}+\mathcal{L}_{\text{perceptual}}(\mathbf{o},\operatorname{dec}(z))$
where $\mathcal{L}_{\text{perceptual}}$ is a perceptual loss (Johnson et al., 2016; Larsen et al., 2016), proposed in (Micheli et al., 2023).
@@
Table 8: The encoder and decoder architectures of the VQ-VAE model. “Conv(a,b,c)” represents a convolutional layer with kernel size $a\times a$, stride of $b$ and padding $c$. A value of $c=\text{Asym.}$. represents an asymmetric padding where a padding of 1 is added only to the right and bottom ends of the image tensor. “GN” represents a GroupNorm operator with $8$ groups, $\epsilon=1e-6$ and learnable per-channel affine parameters. SiLU is the Sigmoid Linear Unit activation (Hendrycks & Gimpel, 2017; Ramachandran et al., 2018). “Interpolate” uses PyTorch’s interpolate method with scale factor of 2 and the “nearest-exact” mode.
@@ A.3.1 Optimization
For each training example in the form of a trajectory segment in token representation $\tau=\mathbf{z}_{1},\mathbf{a}_{1},\ldots,\mathbf{z}_{H},\mathbf{a}_{H}$, the optimization objective is given by
$\mathcal{L}_{\mathcal{M}}(\theta,\phi,\tau)=\sum_{t=1}^{H}\mathcal{L}_{\text{obs}}(\theta,\mathbf{z}_{t},p_{\theta}(\hat{\mathbf{z}}_{t}|\mathbf{Y}^{\text{u}}_{t}))+\mathcal{L}_{\text{reward}}(\theta,r_{t},\hat{r}_{t})-\log(p_{\theta}({d}_{t}|\mathbf{Y}^{\text{u}}_{t}))+\sum_{i=1}^{N_{\text{ens}}}\mathcal{L}_{\text{obs}}(\phi_{i},\mathbf{z}_{t},p_{\phi_{i}}(\hat{\mathbf{z}}_{t}|\operatorname{sg}(\mathbf{Y}^{\text{u}}_{t})))$

### v3

@@ 3 Experiments
To evaluate sample efficiency, we used benchmarks that measure performance within a fixed, limited environment interaction budget. These benchmarks were also selected to address key research questions: (1) whether the proposed multi-modality approach is effective—both in continuous control settings and in handling multi-modal observations; and (2) whether the integrated components are effective across diverse environments (Section 3.3).
@@ 3.1 Experimental Setup
We evaluate Simulus on three sample-efficiency benchmarks of different observation and action modalities: Atari 100K [28], DeepMind Control Suite (DMC) Proprioception 500K [53], and Craftax-1M [35].
@@
The DeepMind Control Suite (DMC) is a set of continuous control tasks involving multiple agent embodiments ranging from simple single-joint models to complex humanoids. Here, we follow the subset of proprioception tasks used for the evaluation of DreamerV3 [20], where observations and actions are continuous vectors. At each task, the agent’s interaction budget is limited to 500K steps.
@@
On Atari-100K, we compare Simulus against DreamerV3 [20] and several methods restricted to image observations: TWM [44], STORM [60], DIAMOND [3], and REM [11]. On DMC, we compare exclusively with DreamerV3, currently the only planning-free world model method with published results on the 500K proprioception benchmark. On Craftax-1M, we compare against TWM [13], a concurrent work that proposes a Transformer based world model with a focus on the Craftax benchmark, and the baselines reported in the Craftax paper: Random Network Distillation (RND) [10], PPO [48] with a recurrent neural network (PPO-RNN), and Exploration via Elliptical Episodic Bonuses (E3B) [23]. As Craftax is a recent benchmark, there are no other published results in existing world models literature. Following the standard practice in the literature, we exclude planning-based methods [21, 57], as planning is an orthogonal component that operates on any given model, typically incurring significant computational overhead.
@@
For Atari, we report human-normalized scores (HNS), calculated as $\frac{\text{agent\_score}-\text{random\_score}}{\text{human\_score}-\text{random\_score}}$ [39]. Following the protocol of Agarwal et al. [2] and using their toolkit, we report the mean, median, interquantile mean (IQM), and optimality gap metrics with 95% stratified bootstrap confidence intervals. For DMC and Craftax, we report the raw agent returns. We use 5 random seeds per environment. In each experiment, final performance is evaluated using 100 test episodes at the end of training and the mean score is reported.
@@ 3.2 Results
Simulus achieves state-of-the-art performance across all three benchmarks (Figure 1). On Atari 100k, it outperforms all baselines across key metrics (Figure 4). Notably, Simulus is the first planning-free world model to reach human-level IQM and median scores, achieving superhuman performance on 13 out of 26 games (Table 9, Appendix B). Building on REM, these gains are attributed to the integration of the proposed components, demonstrating their combined effectiveness.
Effectiveness in continuous environments
Figure 4 provides compelling evidence that token-based architectures can perform well in continuous domains—even with compact vocabularies: Simulus consistently matches DreamerV3 across most tasks and slightly outperforms it on average.
Effectiveness in environments with multi-modal observations
We evaluate multi-modality performance in Craftax, as it combines an image-like 2D grid map with a vector of features, involving multiple tokenizers ($\mathcal{V}$). Simulus maintains sample-efficiency in this multi-modal environment, outperforming both concurrent world model methods (Figure 1) and all model-free baselines (Figure 5), including exploration-focused algorithms. With 444 tokens per observation arranged into sequences of 147 embeddings, even short trajectories in Craftax contain thousands of tokens, demonstrating Simulus’s efficient handling of long sequences. These findings indicate that the world model ($\mathcal{M}$) and controller ($\mathcal{C}$) maintain strong performance under multi-modal inputs when processed by the proposed modular tokenizer.
Figure 6: Ablations results on the Atari-100K and DeepMind Control Proprioception 500K benchmarks. A subset of 8 games was used for each ablation.
@@ 3.3 Ablation Studies
We ablate the intrinsic rewards, prioritized replay, and regression-as-classification to demonstrate their individual contributions to Simulus’s performance. In each experiment, Simulus is modified by disabling a single component. Due to limited computational resources, we consider a subset of 8 tasks for each of the Atari and DMC benchmarks, and exclude Craftax from this analysis. For Atari 100K, we used games in which significant improvements were observed. For DMC, we chose a subset that includes different embodiments. We defer the specific environment names to Appendix C.
The results are presented in Figure 6. Although all components contributed to Simulus’s final performance, intrinsic rewards were especially crucial for achieving competitive performance in both benchmarks. Interestingly, the Atari 100K results indicate that combining all three components yields a significantly stronger algorithm. These findings also suggest that both prioritized world model replay and regression-as-classification enhance the effectiveness of intrinsic rewards.
More broadly, the results in Figure 6 demonstrate that encouraging the controller to explore regions of high epistemic uncertainty through intrinsic rewards significantly improves its performance in world model agents, even in reward-rich environments. This observation is non-trivial in a sample-efficient setting, where the limited interaction budget makes model-driven exploration particularly costly, as it consumes resources that could otherwise be used for task-related exploration during data collection. The latter type of exploration aims to collect new information about the true reward signal, which defines the task and its success metric. On the other hand, model-driven exploration may guide the controller towards environment regions that are irrelevant to the task at hand.
@@ 4 Related Work
Large-scale token sequence models for multi-modal agent trajectories have been proposed in [33, 34, 43, 47]. Gato [43] and TDM [47] tokenize inputs via predefined transformations, while Unified IO [33, 34] leverages pretrained models. These methods do not learn control through RL but rely on expert data. They also use massive models—with billions of parameters, large vocabularies, and significantly more data and compute than sample-efficient world models. Consequently, it remains unclear whether their design choices would be effective in online, sample-efficient settings with non-stationary and limited data.
@@
Model-based agents that learn policies solely from simulated data generated by a learned world model were introduced by Ha and Schmidhuber [17], followed by the influential Dreamer family [18, 19, 20]. Dreamer jointly optimizes its representation and recurrent world models via a KL divergence between learned prior and posterior distributions, leading to interfering objectives (Appendix B.1) and a complex, monolithic architecture that complicates development and scaling [59]. With the rise of Transformer architectures in language modeling [56, 9], Transformer-based Dreamer variants emerged [60, 44], alongside token-based world models (TBWMs) that treat trajectories as discrete token sequences [37, 11]. However, these methods are limited to visual environments with discrete actions (e.g., Atari 100K), leaving their performance in other modalities uncertain. Recently, DIAMOND [3], a diffusion world model inspired by advances in generative modeling [45], was introduced. While it generates visually compelling outputs, it remains limited to visual domains.
Intrinsically Motivated World Model Agents
Although intrinsic motivation (IM) has been extensively studied [41, 10, 22, 5, 4], its use in world models typically involves an exploration pretraining phase followed by limited task-specific fine-tuning [49, 36, 30]. While combining IM with prioritized replay has shown promise [30], it remains unexplored in standard sample-efficiency settings with external rewards.
Large-Scale Video World Models
Building on recent advances in video generative modeling [25, 7, 8], recent works have introduced large-scale video world models [14, 1, 12, 54], trained offline on extensive pre-collected data to predict future frames. However, these models do not address control learning, particularly RL. While recent efforts aim to bridge this gap [59], they remain confined to visual environments and lack comprehensive empirical evaluation.
@@ 5 Limitations and Future Work
Here, we briefly highlight several limitations of this work. First, although the feature quantization approach for tokenizing continuous vectors showed promise, it leads to excessive sequence lengths. We believe that more efficient solutions can be found for dealing with continuous inputs. Second, due to the scarcity of rich multi-modal RL benchmarks, we could not extensively explore diverse modality combinations in our experiments. Lastly, token-based world model agents remain significantly slower to train than other baselines in sample-efficient RL. Nonetheless, their modular design enables faster policy inference as the controller is independent of the world model.
@@ 6 Conclusions
In this paper, we demonstrated the effectiveness of several underutilized techniques for improving world model agents. A modular multi-modality tokenization framework broadens their applicability across diverse domains, while intrinsic motivation, prioritized world model replay, and regression-as-classification enhance sample efficiency, particularly when combined. These techniques were incorporated into a token-based world model, yielding Simulus. Extensive experiments show that Simulus achieves state-of-the-art performance across diverse benchmarks, including visual, continuous, and symbolic domains. It outperforms all baselines on key metrics in both Atari-100K and the challenging Craftax benchmarks. Ablations further highlight the individual contribution of each component. We hope that the highly modular design of Simulus, along with the released code and model weights, provides a strong foundation for future work.
@@ A.1 Hyperparameters
For the DMC benchmark, we use a lower embedding dimension (Table 3) due to the significantly lower dimensionality of its observations compared to other benchmarks. As in prior work (e.g., DreamerV3), we adopt a smaller model for this setting. Note that the reduced number of Retention heads ensures a consistent head dimensionality (64).
Additionally, we used a limited decay range in DMC (Table 3) as observations effectively represent full MDP states, eliminating the need for long-term memory. By constraining the decay range, we explicitly encode this inductive bias into the model.
In the Craftax benchmark, we reduce the number of layers (Table 3) to lower computational cost. The interaction budget in Craftax is 1M steps, resulting in a particularly expensive training process. Here, we increase the decay range as an inductive bias to encourage long-term memory.
The weighting of intrinsic versus extrinsic rewards (Table 4) varies across benchmarks due to differences in reward structure and scale. For instance, DMC provides dense rewards with typical task scores reaching around 1000, whereas Craftax has extremely sparse rewards, with cumulative scores rarely exceeding 20—even over thousands of steps.
@@
All remaining hyperparameters were tuned empirically or based on REM [11], with minimal impact on training cost and adjusted primarily for performance. Due to limited computational resources, we were unable to conduct extensive tuning, and we believe that further optimization could improve Simulus’s performance.
@@ A.2.1 Image Observations
Image observations are tokenized using a vector-quantized variational auto-encoder (VQ-VAE) [55, 15]. A VQ-VAE comprises a convolutional neural network (CNN) encoder, an embedding table $\mathbf{E}\in\mathbb{R}^{n\times d}$, and a CNN decoder. Here, the size of the embedding table $n$ determines the vocabulary size.
@@
$z=\arg\min_{i}\|\mathbf{h}-\mathbf{E}(i)\|,$
where $\mathbf{E}(i)$ is the $i$-th row of $\mathbf{E}$. To reconstruct the original image, the decoder first maps $\mathbf{z}$ to their embeddings using $\mathbf{E}$. During training, the straight-through estimator [6] is used for backpropagating the learning signal from the decoder to the encoder: $\hat{\mathbf{h}}=\mathbf{h}+\operatorname{sg}(\mathbf{E}_{z}-\mathbf{h})$. The architecture of the encoder and decoder models is presented in Table 7.
@@
$\mathcal{L}(\operatorname{enc},\operatorname{dec},\mathbf{E})=\\ \|\mathbf{o}-\operatorname{dec}(z)\|_{2}^{2}+\|\operatorname{sg}(\operatorname{enc}(\mathbf{o}))-\mathbf{E}(z)\|_{2}^{2}+\|\operatorname{sg}(\mathbf{E}(z)-\operatorname{enc}(\mathbf{o})\|_{2}^{2}+\mathcal{L}_{\text{perceptual}}(\mathbf{o},\operatorname{dec}(z)),$
where $\mathcal{L}_{\text{perceptual}}$ is a perceptual loss [27, 32], proposed in [37].
@@
Table 7: The encoder and decoder architectures of the VQ-VAE model. “Conv(a,b,c)" represents a convolutional layer with kernel size $a\times a$, stride of $b$ and padding $c$. A value of $c=\text{Asym.}$. represents an asymmetric padding where a padding of 1 is added only to the right and bottom ends of the image tensor. “GN" represents a GroupNorm operator with $8$ groups, $\epsilon=1e-6$ and learnable per-channel affine parameters. SiLU is the Sigmoid Linear Unit activation [24, 42]. “Interpolate" uses PyTorch’s interpolate method with scale factor of 2 and the “nearest-exact" mode.
@@ A.3.1 Optimization
For each training example in the form of a trajectory segment in token representation $\tau=\mathbf{z}_{1},\mathbf{z}^{\text{a}}_{1},\ldots,\mathbf{z}_{H},\mathbf{z}^{\text{a}}_{H}$, the optimization objective is given by
$\mathcal{L}_{\mathcal{M}}(\theta,\phi,\tau)=\sum_{t=1}^{H}\mathcal{L}_{\text{obs}}(\theta,\mathbf{z}_{t},p_{\theta}(\hat{\mathbf{z}}_{t}|\mathbf{Y}^{\text{u}}_{t}))+\mathcal{L}_{\text{reward}}(\theta,r_{t},\hat{r}_{t})-\log(p_{\theta}({d}_{t}|\mathbf{Y}^{\text{u}}_{t}))\\ +\sum_{i=1}^{N_{\text{ens}}}\mathcal{L}_{\text{obs}}(\phi_{i},\mathbf{z}_{t},p_{\phi_{i}}(\hat{\mathbf{z}}_{t}|\operatorname{sg}(\mathbf{Y}^{\text{u}}_{t}))),$